OpenStudy (anonymous):

At TelePower Plus, long-distance phone calls to China cost $0.59 for the first minute and $0.25 for each additional minute, plus an additional roaming charge of $2.50. If the total charge of a call to Beijing was $11.84, how many minutes did the call last?

OpenStudy (anonymous):

\[.59+250=3.09\] is the flat fee then if you call the amount of minutes \(x\) the total price was \[3.09+.25x\] which you know is \(11.84\) so you can solve \[3.09+.25x=11.84\] for \(x\)

OpenStudy (anonymous):

may be easier to multiply both sides by \(100\) and solve \[309+25x=1184\] with no decimals

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

yea thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

\[\frac{11.84-2.50-0.59}{0.25}+1=36 \min\]